The length of a rectangular mailing label is 3 cm less than twice the width. The perimeter is 48 cm. Find the dimensions of the label.

Answer by Miceandwine(15)

The length of a rectangular mailing label is 3 cm less than twice the width. The perimeter is 48 cm. Find the dimensions of the label.
Okay, so first let us set this up; remember that this is perimeter so the figures will have to be doubled first
2(x+2x-3)=48
2x+4x-6=48 Now combine like terms.
6x=56 Now divide both sides by 6.
x=9 Hooray! We now have one side of the problem. Now, let's plug this number back into the equation given to see what the other side is.
2(9)-3=15 So the other side is 15 cm. Let's check our math:
15+15+9+9=48
I hope this helped!